American Airlines hedges a £2.5 million receivable by selling pounds forward. If the spot

rate is $1.742/£ and the 90-day forward rate is $1.7158/£, what is American's actual true

cost of hedging if the spot rate at the end of 90-days is $1.73/£?
Answer #1

Amount received from hedging as per the forward rate = 2,500,000* 1.7158 =

$4,289,500

Amount that would have been received otherwise = 2500,000* 1.73 =

$4325000

True cost of hedging = 4325,000 - 4289500 = $35,500